Graph the equation: 2y+x=-3.
Answer by elima(1433)

2y+x=-3.
There are diffeerent ways to solve this. One way is to choose some easy points for x, (say 0, 1, and 2), plug them in, and calculate the corresponding y-values.
2y+(0)=-3
2y=-3
y=-3/2
First point; (0,-3/2)
2y+(1)=-3
2y=-3-1
2y=-4
y=-2
2nd point;
(1,-2)
2y+(2)=-3
2y=-3-2
2y=-5
y=-5/2
3rd point;
(2,-5/2)
Now plot these points on a graph, and you have your line.
hope you understand
=)